kristiannnnnt:

A number is chosen at random from the first 100 positive integers. Find the probability that the number is either even or a perfect square. 11/20 1/20 3/5

snowflake0531:

From 1 to 100 there are 50 even numbers And about perfect squares, we have from 1 to 10 \[1^2=1~2^2=4~3^2=9~4^2=16~5^2=25~6^2=36~7^2=49~8^2=64~9^2=81~10^2=100\] so we have 1, 4, 9, 16, 25, 36, 49, 64, 81, 100 And out of these, 4, 16, 36, 64, 100 are even numbers So besides for those 5, there are 5 more that are odd, so there are 55 numbers that are either even or a perfect square So that's 55/100

snowflake0531:

So then from there you can simplify 55/100

kristiannnnnt:

so how you divided ?

snowflake0531:

? I didn't divide anything

kristiannnnnt:

the other guy